# {adjustwidth}0cm-0.2cm Higgs-Induced Gravitational Waves: the Interplay of Non-Minimal Couplings, Kination and Top Quark Mass

### v1

@@ Appendix A Summary of parametric formulas
with $\nu=\sqrt{3\xi/2}$ and $n=-\log_{10}(\lambda)$. The time at which the Higgs energy density achieves a radiation-like scaling is
@@
$\gamma_{1}=33.63+15.02\,n-0.22\,n^{2}\,,\;\gamma_{2}=7.91-0.01\,n+0.02\,n^{2}\,,\;$ (30)
and the Higgs energy density at that timescale is
@@
Appendix B Setting up classical lattice simulations
Our numerical analysis of the Higgs dynamics in its effective potential is carried out using the $\mathcal{C}osmo\mathcal{L}attice$ code Figueroa:2021yhd ; Figueroa:2016wxr following the general procedure introduced in previous works Laverda:2023uqv ; Laverda:2024qjt ; Bettoni:2024ixe . The first step consists in implementing the full effective potential in a model file that includes a curvature-dependent effective mass, the three-loop RGI running of $\lambda(\mu)$ in (3) with an explicit dependence on the top quark and Higgs masses and the additional higher-order operator in (24). The lattice parameters are chosen to guarantee the stability of the solution and the covering of all relevant scales.
  • The number of lattice points per dimension has been fixed to $N=960$ with IR and UV momenta being defined by $\kappa_{\rm IR}=2\pi/L$, $\kappa_{\rm UV}=\sqrt{3}N\kappa_{\rm IR}/2$, where the lattice size is $L=N\,\delta x$ and $\delta x=4\pi\,\nu/N$ is the dimension of a single lattice cell. From the discussion in Bettoni:2019dcw , we set the smallest momentum in the tachyonic band with $\kappa_{\text{IR}}=\mathcal{H}_{\rm kin}$, while the largest amplified momentum is set to be much smaller than the lattice’s UV momentum, i.e. $\sqrt{4\nu^{2}-1}\mathcal{H}_{\rm kin}\ll\kappa_{\text{UV}}$.

### v2

@@ Appendix B Summary of parametric formulas from lattice simulations
    with $\nu=\sqrt{3\xi/2}$ and $n=-\log_{10}(\lambda)$.
  • In the same lattice-based analysis, it is possible to derive the time at which the Higgs energy density achieves a radiation-like scaling $z_{\rm rad}$ and evaluate its energy density at that point. We found
@@
    $\gamma_{1}=33.63+15.02\,n-0.22\,n^{2}\,,\;\gamma_{2}=7.91-0.01\,n+0.02\,n^{2}\,.\;$ (49)
    The Higgs energy density at that timescale is given by
@@
Appendix C Numerical lattice simulations
Our numerical analysis of the Higgs dynamics in its effective potential is carried out using the $\mathcal{C}osmo\mathcal{L}attice$ code Figueroa:2021yhd ; Figueroa:2016wxr following the general procedure introduced in previous works Laverda:2023uqv ; Laverda:2024qjt ; Bettoni:2024ixe . The first step consists in implementing the full effective potential in a model file that includes a curvature-dependent effective mass, the three-loop RGI running of $\lambda(\mu)$ in (10) with an explicit dependence on the top-quark mass and the additional higher-order operator in (31). The lattice parameters are chosen to guarantee the stability of the solution and the covering of all relevant scales.
  • The number of lattice points per dimension has been fixed to $N=960$ for the simulations shown in Figure 7, with IR and UV momenta being defined by $\kappa_{\rm IR}=2\pi/L$, $\kappa_{\rm UV}=\sqrt{3}N\kappa_{\rm IR}/2$, with $\kappa\equiv k/\sqrt{6\xi}\mathcal{H}_{\rm kin}$. The lattice size is $L=N\,\delta x$ and $\delta x=4\pi\,\nu/N$ is the dimension of a single lattice cell. From the discussion in Bettoni:2019dcw , we set the smallest momentum in the tachyonic band with $\kappa_{\text{IR}}=\mathcal{H}_{\rm kin}$, while the largest amplified momentum is set to be much smaller than the lattice’s UV momentum, i.e. $\sqrt{4\nu^{2}-1}\mathcal{H}_{\rm kin}\ll\kappa_{\text{UV}}$.
@@ Appendix C Numerical lattice simulations
Figure 8: Time evolution of the ratio $\Delta_{1}$ for benchmark scenarios with $\lambda=10^{-3}$ and $\nu=10,\;20$ and different scales of kination $\mathcal{H}_{\rm kin}$.
Figure 9: Time evolution of the ratio $\Delta_{2}$ for benchmark scenarios with $\lambda=10^{-3}$ and $\nu=10,\;20$ and different scales of kination $\mathcal{H}_{\rm kin}$.
Following the discussion in Section 2 and in the previous Appendix B, we can asses the spectator field-regime thanks to some benchmark 3+1 lattice simulations. In particular, since we are interested in checking some general properties of the symmetry breaking, it is more instructive to fix the Higgs self-coupling to some constant value. Within the parameter space in Figure 2, the scale-dependence of $\lambda$ leads to typical values $\lambda>10^{-3}$. Therefore, we perform a few lattice simulations with $\lambda=10^{-3}$, $\nu=10,\;20$ and $N=256$ to probe the regime in which the spectator-field approximation is valid. Figure 8 shows the time-evolution of the term $\Delta_{1}$ for different scales of kination. Crossing the value $\Delta_{1}=1$ implies a change in the sign of the Ricci scalar and the restoration of the unbroken phase in the effective potential. This happens generally when the tachyonic mass is large, i.e. for large scales of kination and large non-minimal couplings. Within the range $\mathcal{H}_{\rm kin}\lesssim 10^{12}\textrm{ GeV}$, the energy density of the spectator field is sufficiently smaller than the inflaton energy density not to lead to a faster symmetry restoration.
The second consistency check of the spectator approximation is the smallness of $\Delta_{2}$, which is implied by imposing the cut-off scale of the effective theory on the Higgs fluctuations. In this case, we can make use of the parametric formula for $\langle h_{\rm br}^{2}\rangle$ to estimate the constraint $\langle h_{\rm br}^{2}\rangle\ll M_{P}^{2}/\xi^{2}$ in a parametric way. The result is already included in figure 2 and figure 4 as a light-blue-shaded area and it takes into account the running of the Higgs self-coupling. We maintained a conservative bound of $\langle h_{\rm br}^{2}\rangle=10^{-2}\times M_{P}^{2}/\xi^{2}$, thus accounting for the small discrepancy between the average amplitude at backreaction time and the largest average amplitude at the end of the first semi-oscillation. Figure 9 shows the full evolution of $\Delta_{2}$ during the tachyonic phase and the beginning of the non-linear phase for a small number of benchmark points in our parameter space. Notice again, that the value of $\lambda=10^{-3}$ is chosen in a conservative way and that the typical value of the Higgs self-coupling in the vast majority of our parameter space is larger. The $1\%$ threshold excludes kination scales above $\sim 10^{12}\textrm{ GeV}$, which we have taken as an upper limit for Figures 3, 5, 6.